# Surprisingly Strong Performance Prediction with Neural Graph Features

### v1

@@ 3 Introducing Neural Graph Features
Figure 1: ZCP score of all networks against the validation accuracy - colors indicate the number of conv3x3 and 1x1
@@ 3.1 Limitations of zero-cost proxies
Figure 2: Spearman correlation of zero-cost proxies with validation accuracy in clusters of architectures with the same number of 1x1 and 3x3 convolutions (a,b), and only 3x3 convolutions (c). 0-6 represents clusters with corresponding convolution counts, all data is the correlation over all available networks (same as in NB-Suite-Zero (Krishnakumar et al., 2022)).
@@ 3.2 Neural graph features
Figure 4: Validation accuracy prediction across 3 benchmarks with 1024 sampled networks. Comparison of 3 encodings (ZCP, onehot, and GRAF) and their combinations. Blue boxes denote runs including GRAF.
@@ 3.3 Using Graph Features in Prediction
Figure 4 shows results for 1024 sampled train networks on cifar-10 and NB101, NB201, and NB301. GRAF performs better than ZCP and onehot in all cases, and ZCP + GRAF yields the overall best results. On smaller sample sizes (32, 128), GRAF tends to be similar or only slightly worse than ZCP (Figure 7), but ZCP + GRAF outperform ZCP only. For TNB-micro (Figures 15, 16) and TNB-macro (Figures 17, 18), GRAF is better than ZCP + onehot on most datasets and sample sizes, and sometimes even better than ZCP + GRAF.
These results show that while zero-cost proxies have some flaws, they still capture important network properties that help the prediction. The strong performance of ZCP + GRAF suggests that GRAF does not capture all network properties, and ZCP and GRAF complement each other well for prediction.
@@ 3.4 Feature importance
Table 1: The ten most important features on NB201 cifar-10 and TNB101-micro - autoencoder from ranking based on Shapley values computed from random forest trained on 1024 training samples with ZCP+GRAF.
@@ 3.5 Analysing feature redundancy
First, we analyze the contribution of the following feature groups: max path length (MAX), min path length (MIN), operation count (OPS), and node degree (DEG). We look at performance of ZCP + one feature group, and ZCP + all but one feature group. We compare two cases: with all ZCP, and only with flops and params (FP). Figure 5 shows random forest prediction across 50 seeds on 1024 networks (NB201, cifar-10). We see that DEG has the highest individual contribution. However, when using all but one group, leaving out MIN leads to the greatest drop in performance. It also seems that OP do not have a large influence when other feature groups are present, indicating they might possibly be computed from the other features or from FP. However, we still use them in our models for better interpretability. Figures 19 and 20 in the appendix show similar results for NB101 and NB201 + ImageNet16-120 — all groups except OP are needed, but the most influential groups differ.
@@ 4 Empirical Evaluation on Diverse Tasks
Figure 7: Robust accuracy prediction for the PGD adversarial attack with 1024 training data on the attack perturbations $\epsilon=1.0/255$ for the objectives, only robust accuracy (top), and both, validation and robust accuracy (bottom).
@@ 4.1 Other Prediction Tasks
In the following section, we evaluate how GRAF improves the results on the task of predicting the robustness of the architectures. We use the evaluated robustness of the NAS-Bench-201 (Dong & Yang, 2020b) from the robustness dataset (Jung et al., 2023). This dataset evaluated the architectures from NAS-Bench-201 on different adversarial attacks and different perturbation strength; we consider in this paper two white-box attacks (FGSM (Goodfellow et al., 2015), and PGD (Kurakin et al., 2017). Important to note that the adversarial robustness is highly correlated with the validation accuracy since the former can never improve over the latter. Here, we will differentiate the robustness evaluation into two different tasks: (i) predicting only the robustness accuracy, and (ii) predicting both, the validation accuracy and the robustness accuracy jointly (Multi_Obj). In more detail, we evaluate the prediction ability on both attacks for three different perturbation strengths ($\epsilon=\{0.1/255,0.5/255,1.0/255\}$) on cifar-10 using three different training sizes for learning the random forest prediction model. For the first prediction type (i), the combination of zero-cost proxies, the proposed graph features, and the additional onehot encoding shows the highest Kendall tau in all considered tasks (see Section K in the appendix for a detailed overview). Furthermore, with increasing perturbation strength, using only the graph features shows a better prediction ability than using only the zero-cost proxies from the literature. The same behavior is visible for the joint objective prediction task (ii). The latter task of predicting both accuracies (validation and robust) at once shows also a higher Kendall tau. Thus, this seems to be the easier task for the prediction model, which was also shown in Lukasik et al. (2023). We present the boxplots for the PGD attack for perturbation strength $\epsilon=1.0/255$ for both evaluation tasks in 7 and more results on both adversarial attacks in both settings in Section K in the appendix.
@@ 4.2 Other Types of Encoding
Apart from the GRAF features and the zero-cost proxies, we also experimented with other types of network architecture encodings – path encoding, arch2vec (Dong & Yang, 2020a), and NASBOWL’s Weisfeiler-Lehman features (WL) for the number of iterations equal to $1$. (Ru et al., 2021). Detailed results of these experiments are included in the appendix (Tables 19 and 20). Here, we give just a brief overview.
@@ 4.3 Existing performance predictors
On the accuracy task (cf. Table 27) on NB201 BRP-NAS is slightly better than the RF-based models only with the largest training size for the ImageNet16-120 target. On the NB301 benchmark, however, BRP-NAS has the same or better results for all training sizes, which may indicate that the GCN may be able to extract additional important information from the graph structure. On the hardware tasks (Table 28) , the original BRP-NAS has a similar performance to the best combination of features with RF-based predictors. Including the other information improves performance in all cases, with BRP-NAS + GRAF + OH being among the best configurations. Interestingly, very often the performance of BRP-NAS + ZCP is not much better than the performance of BRP-NAS alone. On the other hand, on the robustness tasks (Tables 29,30), BRP-NAS has slightly worse results than RF-based models with GRAF and onehot features. Additionally, unlike the other tasks, additional information does not improve the results, except for the smallest perturbation and the smallest training set. An explanation could be that the properties relevant for prediction are captured both by BRP-NAS and ZCP + GRAF.
@@
Figure 8: Evaluation of NASLib performance predictors on a subset of NB101, cifar-10
@@ 4.4 NASLib search run
Figure 9: Bayesian optimization – search on ImageNet16-120
@@ 6 Discussion
7 Broader Impact
@@ C.2 Unreachable Branches in NB201 and TNB101-micro
All in all, we believe removing unreachable operations in NAS experiments is good practice – these networks should not be used in practice anyway (due to higher energy costs), and removing the operations requires just a DFS run, i.e. $\mathcal{O(V+E)}$ for $V$ nodes and $E$ edges.
@@ Appendix H Feature Redundancy
We list additional plots of the feature group ablation in Figures 19 and 20.
@@ Appendix I Other Encoding Types
In this section, we list the full results of other encoding types (listed in Section C.4) in Tables 19 and 20. It is important to note that NASBOWL uses the WL features in a Gaussian process predictor, and their performance in a random forest might be limited.
@@
Figure 21 shows the results for edgegpu_energy prediction. The following Tables 21, 22 and 23 list average Kendall tau for selected nontrivial tasks and various predictors. In all cases, the best results are produced with settings containing GRAF among predictors. The similar results were obtained also on the rest of the tasks from HW-NAS-Bench, the only exceptions are 5 tasks, where onehot + FP produced the best result (from total 90 tasks; three datasets, three different training sizes, 10 hardware metrics).
@@
TA-GATES requires over 10 minutes to train on a single seed (for all sample sizes), but ZCP + GRAF + XGB+ needs only 12 seconds on the largest sample size. The total time required to run all 50 repetitions and 4 sample sizes is listed in Table 31 (on Intel Xeon CPU E5-2620, NVIDIA GeForce GTX 1080 Ti, no networks from the search space were trained since we use precomputed search space data). Figure 24 shows the results of all models and shows that ZCP + GRAF is competitive with TA-GATES at a fraction of the train time.
Table 31: Runtimes of different models across all experiments. ZCP + GRAF experiments take less then 1 hour on a CPU, while TA-GATES needs over 37 GPU hours.
@@
Figure 25: NASLib predictor evaluation, NAS-Bench-101 and cifar-10
Figure 26: NASLib predictor evaluation, NAS-Bench-201 and cifar-10
@@
Figure 28: NASLib predictor evaluation, NAS-Bench-301 and cifar-10
@@
Figures 29 and 30 show results of the search on cifar-10 and ImageNet16-120 respectively.
Figure 29: NASLib search experiment, NAS-Bench-201 and cifar-10
Figure 30: NASLib search experiment, NAS-Bench-201 and ImageNet16-120

### v2

@@ 1 Introduction
We make the code of our contributions available publicly [^1] https://github.com/gabikadlecova/zc_combine
@@ 3 Introducing Neural Graph Features
Figure 1: ZCP score of all networks from NB201 against the validation accuracy - colors indicate the number of conv3x3 and 1x1
@@ 3.1 Limitations of zero-cost proxies
Figure 2: Spearman correlation of zero-cost proxies with validation accuracy by clusters of architectures with the same number of 1x1 and 3x3 convolutions (a,b), and only 3x3 convolutions (c). 0-6 represents clusters with corresponding convolution counts, all data is the correlation over all available networks (same as in NB-Suite-Zero (Krishnakumar et al., 2022)).
@@ 3.2 Neural graph features
Figure 4: Validation accuracy prediction (cifar-10) across 3 benchmarks with 1024 sampled networks. Comparison of 3 encodings (ZCP, onehot, and GRAF) and their combinations. Blue boxes denote runs including GRAF.
@@ 3.3 Using Graph Features in Prediction
Figure 4 shows results for 1024 sampled train networks on cifar-10 and NB101, NB201, and NB301. GRAF performs better than ZCP and onehot in all cases, and ZCP + GRAF yields the overall best results. Figure 5 shows results for the autoencoder task and TNB101-micro. We can see that using GRAF leads to better results than ZCP or OH across all sample sizes. For other TNB-micro tasks (Figures 16, 17) and TNB-macro (Figures 18, 19), the best configurations always include GRAF. Overall, GRAF performs the same or better than ZCP even on the smaller sample sizes (32, 128) across all benchmarks and tasks (Section F in the appendix) except for NB201, where it is slightly worse (Table 7).
These results show that while zero-cost proxies have some flaws, they still capture important network properties that help the prediction. The strong performance of ZCP + GRAF suggests that GRAF does not capture all network properties, and ZCP and GRAF complement each other well for prediction. It is important to note that for some tasks beyond image classification, using GRAF on its own is sufficient.
@@ 3.4 Feature importance
Table 1: The ten most important features on NB201 cifar-10 and TNB101-micro - autoencoder. Ranking is based on Shapley values computed from random forest trained on 1024 training samples with ZCP+GRAF.
@@
Figure 5: TNB101_micro autoencoder task across 3 train sample sizes. Comparison of 3 encodings (ZCP, onehot, and GRAF) and their combinations. Blue boxes denote runs including GRAF.
@@ 3.5 Analysing feature redundancy
First, we analyze the contribution of the following feature groups: max path length (MAX), min path length (MIN), operation count (OPS), and node degree (DEG). We look at performance of ZCP + one feature group, and ZCP + all but one feature group. We compare two cases: with all ZCP, and only with flops and params (FP). Figure 6 shows random forest prediction across 50 seeds on 1024 networks (NB201, cifar-10). We see that DEG has the highest individual contribution. However, when using all but one group, leaving out MIN leads to the greatest drop in performance. It also seems that OP do not have a large influence when other feature groups are present, indicating they might possibly be computed from the other features or from FP. However, we still use them in our models for better interpretability. Figures 20 and 21 in the appendix show similar results for NB101 and NB201 + ImageNet16-120 — all groups except OP are needed, but the most influential groups differ.
@@ 4 Empirical Evaluation on Diverse Tasks
Figure 8: Robust accuracy prediction for the PGD adversarial attack with 1024 training data on the attack perturbations $\epsilon=1.0/255$ for the objectives, only robust accuracy (top), and both, validation and robust accuracy (bottom). NB201 and cifar-10.
@@ 4.1 Other Prediction Tasks
In the following section, we evaluate how GRAF improves the results on the task of predicting the robustness of the architectures. We use the evaluated robustness of the NAS-Bench-201 (Dong & Yang, 2020b) from the robustness dataset (Jung et al., 2023). This dataset evaluated the architectures from NAS-Bench-201 on different adversarial attacks and different perturbation strength; we consider in this paper two white-box attacks (FGSM (Goodfellow et al., 2015), and PGD (Kurakin et al., 2017). Important to note that the adversarial robustness is highly correlated with the validation accuracy since the former can never improve over the latter. Here, we will differentiate the robustness evaluation into two different tasks: (i) predicting only the robustness accuracy, and (ii) predicting both, the validation accuracy and the robustness accuracy jointly (Multi_Obj). In more detail, we evaluate the prediction ability on both attacks for three different perturbation strengths ($\epsilon=\{0.1/255,0.5/255,1.0/255\}$) on cifar-10 using three different training sizes for learning the random forest prediction model. For the first prediction type (i), the combination of zero-cost proxies, the proposed graph features, and the additional onehot encoding shows the highest Kendall tau in all considered tasks (see Section K in the appendix for a detailed overview). Furthermore, with increasing perturbation strength, using only the graph features shows a better prediction ability than using only the zero-cost proxies from the literature. The same behavior is visible for the joint objective prediction task (ii). The latter task of predicting both accuracies (validation and robust) at once shows also a higher Kendall tau. Thus, this seems to be the easier task for the prediction model, which was also shown in Lukasik et al. (2023). We present the boxplots for the PGD attack for perturbation strength $\epsilon=1.0/255$ for both evaluation tasks in 8 and more results on both adversarial attacks in both settings in Section K in the appendix.
@@ 4.2 Other Types of Encoding
Apart from the GRAF features and the zero-cost proxies, we also experimented with other types of network architecture encodings – path encoding, arch2vec (Dong & Yang, 2020a), and NASBOWL’s Weisfeiler-Lehman features (WL) for the number of iterations equal to $1$. (Ru et al., 2021). Detailed results of these experiments are included in the appendix (Tables 20 and 21). Here, we give just a brief overview.
@@ 4.3 Existing performance predictors
On the accuracy task (cf. Table 28) on NB201 BRP-NAS is slightly better than the RF-based models only with the largest training size for the ImageNet16-120 target. On the NB301 benchmark, however, BRP-NAS has the same or better results for all training sizes, which may indicate that the GCN may be able to extract additional important information from the graph structure. On the hardware tasks (Table 29) , the original BRP-NAS has a similar performance to the best combination of features with RF-based predictors. Including the other information improves performance in all cases, with BRP-NAS + GRAF + OH being among the best configurations. Interestingly, very often the performance of BRP-NAS + ZCP is not much better than the performance of BRP-NAS alone. On the other hand, on the robustness tasks (Tables 30,31), BRP-NAS has slightly worse results than RF-based models with GRAF and onehot features. Additionally, unlike the other tasks, additional information does not improve the results, except for the smallest perturbation and the smallest training set. An explanation could be that the properties relevant for prediction are captured both by BRP-NAS and ZCP + GRAF.
@@
Figure 9: Evaluation of NASLib performance predictors on a subset of NB101, cifar-10.
@@ 4.4 NASLib search run
Figure 10: Bayesian optimization – search on ImageNet16-120 and NB201.
@@ 6 Discussion
Acknowledgements
This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828. Computational resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic. This research was partially supported by the SVV project number 260 698. G. Kadlecová was supported by Charles University Grant Agency project no. 246322. P. Vidnerová was supported by the project 22-02067S of the Czech Science Foundation. J. Lukasik acknowledges support by the DFG research unit 5336 - Learning to Sense. Frank Hutter is a Hector Endowed Fellow at the ELLIS Institute Tübingen. We would like to thank Arjun Krishnakumar for the valuable discussions that helped to shape the initial ideas of this paper.
Impact Statement
@@ C.2 Unreachable Branches in NB201 and TNB101-micro
All in all, we believe removing unreachable operations in NAS experiments is good practice – these networks should not be used in practice (due to higher energy costs), and removing the operations requires just a DFS run, i.e. $\mathcal{O(V+E)}$ for $V$ nodes and $E$ edges. In related work, NAS-Bench-ASR includes removing unreachable branches in the same manner as we do for NB201 (Mehrotra et al., 2021).
@@ Appendix H Feature Redundancy
We include an additional experiment in Table 19 – we study how the results differ when GRAF is used only with data-dependent proxies (as per (Krishnakumar et al., 2022)). They seem to capture most of the performance gain, and including params leads to the same performance as with all ZCP.
Table 19: Results of GRAF used along with data-dependent ZCP (DD-ZCP), params (P) and all ZCP (ZCP).
benchmark nb101 nb201 nb301
train size 32 128 1024 32 128 1024 32 128 1024
features
GRAF + DD_ZCP $0.56^{0.07}$ $0.66^{0.04}$ $0.73^{0.01}$ $0.69^{0.03}$ $0.78^{0.01}$ $0.84^{0.0}$ $0.4^{0.05}$ $0.53^{0.03}$ $0.63^{0.01}$
GRAF + DD_ZCP + P $0.55^{0.07}$ $0.66^{0.04}$ $0.73^{0.01}$ $0.7^{0.04}$ $0.78^{0.01}$ $0.84^{0.0}$ $0.4^{0.05}$ $0.53^{0.03}$ $0.64^{0.01}$
We list additional plots of the feature group ablation in Figures 20 and 21.
@@ Appendix I Other Encoding Types
In this section, we list the full results of other encoding types (listed in Section C.4) in Tables 20 and 21. It is important to note that NASBOWL uses the WL features in a Gaussian process predictor, and their performance in a random forest might be limited.
@@
Figure 22 shows the results for edgegpu_energy prediction. The following Tables 22, 23 and 24 list average Kendall tau for selected nontrivial tasks and various predictors. In all cases, the best results are produced with settings containing GRAF among predictors. The similar results were obtained also on the rest of the tasks from HW-NAS-Bench, the only exceptions are 5 tasks, where onehot + FP produced the best result (from total 90 tasks; three datasets, three different training sizes, 10 hardware metrics).
@@
TA-GATES requires over 10 minutes to train on a single seed (for all sample sizes), but ZCP + GRAF + XGB+ needs only 12 seconds on the largest sample size. The total time required to run all 50 repetitions and 4 sample sizes is listed in Table 32 (on Intel Xeon CPU E5-2620, NVIDIA GeForce GTX 1080 Ti, no networks from the search space were trained since we use precomputed search space data). ZCP computation time was not included for both models, as TA-GATES provides only precomputed scores. Figure 25 shows the results of all models and shows that ZCP + GRAF is competitive with TA-GATES at a fraction of the train time.
Table 32: Runtimes of different models across all experiments. ZCP + GRAF experiments take less then 1 hour on a CPU, while TA-GATES needs over 37 GPU hours.. Time for computing ZCP is not included.
@@
Figure 26: NASLib predictor evaluation, NAS-Bench-101 and cifar-10.
Figure 27: NASLib predictor evaluation, NAS-Bench-201 and cifar-10.
@@
Figure 29: NASLib predictor evaluation, NAS-Bench-301 and cifar-10.
@@
Figures 30 and 31 show results of the search on cifar-10 and ImageNet16-120 respectively.
Figure 30: NASLib search experiment, NAS-Bench-201 and cifar-10.
Figure 31: NASLib search experiment, NAS-Bench-201 and ImageNet16-120.